OpenStudy (anonymous):

Use the arc length formula and the given information to find s. Show your work for full credit. r = 15 ft θ = 35° s = ?

ganeshie8 (ganeshie8):

first convert the angle to more sophisticated units : radians

ganeshie8 (ganeshie8):

know how to convert \(35\) degrees to radians ?

OpenStudy (anonymous):

yes is it \[\Pi/12\]

OpenStudy (anonymous):

oh nvm i think i did the wrong number

OpenStudy (anonymous):

i did 15 rather than 35

OpenStudy (anonymous):

so im confused on how to convert 35 degrees to radians because i get a repeating decimal answer

OpenStudy (anonymous):

how do i do it @ganeshie8

OpenStudy (anonymous):

\[x^ \circ \times \frac{ \pi }{ 180^\circ }\] to convert into radians

OpenStudy (anonymous):

where x is the given degree

OpenStudy (anonymous):

i get a decimal answer on the bottom though so i dont know where to go from there @iambatman

OpenStudy (anonymous):

because 35 isnt a factor of 180

OpenStudy (anonymous):

\[7\Pi/36\]?

ganeshie8 (ganeshie8):

looks good!

OpenStudy (anonymous):

That's fine now use \[s=r \theta \]

OpenStudy (anonymous):

i got \[105\Pi/36\]

OpenStudy (anonymous):

it doesnt look right

ganeshie8 (ganeshie8):

\[\large s = r*\theta = 15*\dfrac{35\Pi}{180} = \dfrac{35\Pi}{12}\]

ganeshie8 (ganeshie8):

your answer is also correct, its just its not fully simplified